# LapSum – One Method to Differentiate Them All: Ranking, Sorting and Top-k Selection

### v1

@@ Appendix D Experiments
We adopt the identical experimental setup as (Petersen et al., 2022b) using the Adam optimizer (Kingma & Ba, 2015) for all models. A grid search was conducted to fine-tune both the $\alpha$ parameter of $\mathrm{LapSum}$ and the learning rates. We select $\alpha$ from {-1.5, -1, -0.5, -0.2, -0.1} and Leaning Rate (LR) from {$10^{-4.75}$, $10^{-4.5}$, $10^{-4.25}$, $10^{-4}$, $10^{-3.75}$, $10^{-3.5}$, $10^{-3.25}$, $10^{-3}$}. We use early stopping approach. See  Fig. 8 for influence of $\alpha$ parameter on model performance. Baselines As baselines, we reuse scores reported in (Petersen et al., 2022b). Implementation We use code from https://github.com/Felix-Petersen/difftopk kindly made available by Petersen et al. (2022b). Hyperparameter details are available in  Tab. 5
@@
We adopt an identical experimental setup as (Xie et al., 2020), which follows the approach of Grover et al. (2019). A grid search was conducted to fine-tune both the $\alpha$ parameter of $\mathrm{LapSum}$ and the learning rates. We set $\alpha$ as network trainable parameter. We select $\alpha$ at $t_{0}$ from {-10, -5, -3, -1.5, -1, -0.5, -0.2, -0.1} and Leaning Rate (LR) from {$10^{-4.75}$, $10^{-4.5}$, $10^{-4.25}$, $10^{-4}$, $10^{-3.75}$, $10^{-3.5}$, $10^{-3.25}$, $10^{-3}$, $10^{-2.75}$}. We use early stopping approach. For details see Tab. 6. Baselines The baseline results are taken directly from Xie et al. (2020), which in turn references Grover et al. (2019) and He et al. (2016). Implementation We use the code kindly made available by Xie et al. (2020), which itself builds on Cuturi et al. (2019) and Grover et al. (2019).
Table 6: $k$NN experiments hyperparameters.
@@
Figure 9: This figure presents the results from one of the experiments in Sec. 5.2, showing the cost function (blue line) and the learning rate alpha (orange line) over training iterations. The cost function decreases as the model learns, indicating convergence. Parameter $\alpha$ starts at 0.4, initially decreases to allow for finer updates, and then increases to a value above 1 to accelerate learning in later stages.
@@
Figure 10: The figure presents the relationship between the dimension ($n$), shown on the horizontal axis, and the maximum memory usage and computation time, represented on the vertical axis, for several functions with a fixed $k=5$. All calculations displayed in this graph were performed on a CPU. The evaluation examined memory consumption and execution time during the forward pass (left column) and the combined forward and backward passes (right column).
Figure 11: The relationship between the data dimension ($n$), represented on the horizontal axis, and the maximum memory usage and computation time, represented on the vertical axis. The parameter $k$ was dependent on $n$, calculated using the formula $k=n/2$. The computations depicted in this figure were performed on a CPU. The relationships between time and memory usage were analyzed during the forward process (charts in the left column) and the forward and backward processes (charts in the right column).
Figure 12: The figure presents the relationship between the dimension $n$, shown on the horizontal axis, and the maximum memory usage and computation time, represented on the vertical axis, for several functions with a fixed $k=5$. All calculations displayed in this graph were performed on a GPU. The evaluation examined memory consumption and execution time during the forward pass (left column) and the combined forward and backward passes (right column).
Figure 13: The figure illustrates the correlation between the dimensionality ($n$), plotted on the horizontal axis, and the peak memory usage as well as computation time, displayed on the vertical axis, for various functions with $k=n/2$. All computations depicted in this graph were executed on a GPU. The analysis focused on memory utilization and runtime during both the forward pass (left column) and the combined forward and backward passes (right column).

### v2

@@ Appendix D Experiments
We adopt the identical experimental setup as (Petersen et al., 2022b) using the Adam optimizer (Kingma & Ba, 2015) for all models. A grid search was conducted to fine-tune both the $\alpha$ parameter of $\mathrm{LapSum}$ and the learning rates. We select $\alpha$ from {-1.5, -1, -0.5, -0.2, -0.1} and Leaning Rate (LR) from $10^{-4.75}$, $10^{-4.5}$, $10^{-4.25}$, $10^{-4}$, $10^{-3.75}$, $10^{-3.5}$, $10^{-3.25}$, $10^{-3.0}$}, using early stopping approach too. See  Fig. 8 for influence of $\alpha$ parameter on model performance. Baselines As baselines, we reuse scores reported in (Petersen et al., 2022b). Implementation We use code from https://github.com/Felix-Petersen/difftopk kindly made available by Petersen et al. (2022b). Hyperparameter details are available in  Tab. 5.
@@
We adopt an identical experimental setup as (Xie et al., 2020), which follows the approach of Grover et al. (2019). A grid search was conducted to fine-tune both the $\alpha$ parameter of $\mathrm{LapSum}$ and the learning rates. We set $\alpha$ as network trainable parameter. We select $\alpha$ at $t_{0}$ from {-10, -5, -3, -1.5, -1, -0.5, -0.2, -0.1} and Leaning Rate (LR) from $10^{-4.75}$, $10^{-4.5}$, $10^{-4.25}$, $10^{-4}$, $10^{-3.75}$, $10^{-3.5}$, $10^{-3.25}$, $10^{-3.0}$, $10^{-2.75}$}, also using early stopping. For details see Tab. 6. Baselines The baseline results are taken directly from Xie et al. (2020), which in turn references Grover et al. (2019) and He et al. (2016). Implementation We use the code kindly made available by Xie et al. (2020), which itself builds on Cuturi et al. (2019) and Grover et al. (2019).
Table 6: $k$-NN experiments’ hyperparameters.
@@
Figure 9: Results from one of the experiments in Sec. 5.2, showing the cost function (blue line) and the learning rate alpha (orange line) over training iterations. The cost function decreases as the model learns, indicating convergence. Parameter $\alpha$ starts at 0.4, initially decreases to allow for finer updates, and then increases to a value above 1 to accelerate learning in later stages.
@@
Figure 10: Relationship between the dimension ($n$), shown on the horizontal axis, and the maximum memory usage and computation time on vertical axis, for several functions with a fixed $k=5$. All calculations displayed in this graph were performed on a CPU. The evaluation examined memory consumption and execution time during the forward pass (left column) and the combined forward and backward passes (right column).
Figure 11: Relationship between the data dimension ($n$), represented on the horizontal axis, and the maximum memory usage and computation time, represented on the vertical axis. The parameter $k$ was dependent on $n$, calculated using the formula $k=n/2$. The computations depicted in this figure were performed on a CPU. The relationships between time and memory usage were analyzed during the forward process (charts in the left column) and the forward and backward processes (charts in the right column).
Figure 12: Relationship between the dimension $n$ (horizontal axis) and the maximum memory usage and computation time, represented on the vertical axis, for several functions with a fixed $k=5$. All calculations displayed in this graph were performed on a GPU. The evaluation examined memory consumption and execution time during the forward pass (left column) and the combined forward and backward passes (right column).
Figure 13: The figure illustrates the correlation between the dimensionality $n$ on horizontal axis and the peak memory usage as well as computation time, on vertical axis, for various functions with $k=n/2$. All computations depicted in this graph were executed on a GPU. The analysis focused on memory utilization and runtime during both the forward pass (left column) and the combined forward and backward passes (right column).